60. IllI You are driving to the grocery store at 20 m/s. You are 110 m from an intersection when the traffic light turns red. Assume that your reaction time is 0.70 s and that your car brakes with constant acceleration. a. How far are you from the intersection when you begin to apply the brakes? b. What acceleration will bring you to rest right at the intersection? How long does it take you to stop? c.
Answer #1

a)
reaction time = 0.70 s
distance travelled in reaction time = v*t
= 20 m/s * 0.70 s
= 14 m

So, when brake is applied, distance remaining= 110 m - 14 m = 96 m
Answer: 96 m

b)
vf = 0 m/s
d = 96 m
vi = 20 m/s
use:
vf^2 = vi^2 + 2*a*d
0 = 20^2 + 2*a*96
-400 = 2*a*96
a = -2.08 m/s^2
Answer: -2.08 m/s^2

c)
use:
vf = vi + a*t
0 = 20 - 2.08*t
t = 9.6 s

Answer: 9.6 s